A plane takes off from the airport traveling at 200 mph. A jet leave two hours later, traveling at 500mph. How long before the jet overtakes the plane in the air?
Answer by mananth(16949)

A plane takes off from the airport traveling at 200 mph. A jet leave two hours later, traveling at 500mph. How long before the jet overtakes the plane in the air?
plane speed = 200 mph
distance traveled in 2 hours = 400 miles
..
catch up distance = 400 miles
catch up speed = 500mph-200mph=300 mph
..
time = d/s
time = 400/300
time = 4/3 hours
1 hour 20 minutes